OpenStudy (anonymous):

At a school football game, an adult ticket costs $4 and a student ticket costs $1. A total of 250 people attend the game, earning the school $427. How many students attended the game?

OpenStudy (anonymous):

A. 191 B. 59 C. 215 D. 35

OpenStudy (anonymous):

okay

OpenStudy (nurali):

Let A=adults; B=students A+B=250 (total number of people) $4(A)+$1(B)=$427 (total earned) or 4A+B=427 A+B=250 B=250-A 4A+B=427 substitute B with (250-A) 4A+250-A=427 4A-A+250=427 3A+250=427 3A=427-250 3A=177 A=177 / 3 A=59 adults A+B=250 B=250-A B=250-59 B=191 students the answer is A) 191